OpenStudy (adalrymple):

Match each system of linear equations with the quadrant in which the solution lies.

OpenStudy (adalrymple):

-2x + y = 10 x + y = 2 2x - 5y = 10 x + 2y = -14 2x + 4y = -16 3x + y = 9 -3x + y = 9 5x + y = 15